OpenStudy (anonymous):

The equation of the line that goes through the points (-8,7) and (3,3) can be written in general form Ax+Bx+C=0 where: A= B= C=

hero (hero):

You can at least begin with y = mx + b and insert points this way.

OpenStudy (anonymous):

I did & I came up with y=-4/11x+45/11

hero (hero):

you began with 7 = -8m + b 3 = 3m + b ?

hero (hero):

@RoLexx

OpenStudy (anonymous):

why did you start like this (I'm curious & will like to know)

hero (hero):

Because if you did that, you could find m and b this way 7 = -8m + b 3 = 3m + b 7 + 8m = b 3 - 3m = b b = b 7 + 8m = 3 - 3m 8m + 3m = 3 - 7 11m = -4 m = -4/11 b = 7 + 8(-4/11) b = 7 + -32/11 b = 77/11 - 32/11 b = 45/11 y = -4/11x + 45/11 y = (-4x + 45)/11 11y = -4x + 45 4x + 11y = 45 4x + 11y - 45 = 0

hero (hero):

@RoLexx

hero (hero):

therefore A = 4 B = 11 C = -45

OpenStudy (anonymous):

OMG That makes so much sense thanks @Hero